Suppose that a random sample of five observations was taken from a normal population whose variance is 25. The results are 8, 15, 12, 6, 7. Find the 99% confidence interval estimate of the population mean.


1
Expert's answer
2021-03-16T02:40:35-0400

"n=5, \sigma^2=25"

Sample mean


"\bar{X}=\dfrac{8+15+12+6+7}{5}=9.6"

Standard deviation


"\sigma=\sqrt{\sigma^2}=\sqrt{25}=5"

The critical value for "\alpha=0.01" is "z_c=z_{1-\alpha/2}=2.576."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-z_c\times \dfrac{\sigma}{\sqrt{n}},\bar{X}+z_c\times \dfrac{\sigma}{\sqrt{n}})"

"=(9.6-2.576\times \dfrac{5}{\sqrt{5}},9.6+2.576\times \dfrac{5}{\sqrt{5}})"

"=(3.84,15.36)"